OpenStudy (anonymous):

What is the sum of the first 11 terms of the arithmetic series in which a1 = 12 and d = 5? A 38 B 549 C 143 D 209

OpenStudy (anonymous):

sum S = (n/2) × (2a1 + (n-1)d) calculate S by putting n=11

OpenStudy (anonymous):

so putting value in the above formula it is cumming 407 and that is not any of the options

OpenStudy (anonymous):

if that so options may be wrong.

OpenStudy (anonymous):

sum (1 to n) = (1/2)(n)(n+1) 5 sum (1 to n) = (5) (1/2)(n)(n+1) = 5(11)(12)/2 = 330 11 a1= (11)12= 132 I could not get any of the options by brute force 12 + 17 + 22 +...